# A data augmentation strategy for deep neural networks with application to epidemic modelling

### v1

@@ 5 Conclusion
In this study, we explored the impact of a data augmentation strategy on enhancing the forecasting capabilities of Feed-Forward Neural Networks (FNNs) and Nonlinear Autoregressive Networks (NARs) in analyzing and predicting the dynamics of the COVID-19 epidemic during the post-lockdown phase in Italy and Spain. By integrating real data with synthetic data generated from a generalized SIR model, we demonstrated the effectiveness of neural networks in capturing complex epidemic dynamics. FNNs performed well in interpolating within observed data, while NARs were particularly effective in capturing temporal dependencies, making them suitable for long-term forecasting. The generalized SIR model provided a robust foundation for generating high-resolution synthetic data, enriching the neural network training process. The key outcome of this work is the development of a hybrid methodology that bridges the gap between traditional mathematical modeling and machine learning techniques, such as Physics-Informed Neural Networks (PINNs), offering a scalable and flexible framework for epidemic modeling and prediction in real-world scenarios.
@@
Acknowledgments
This work has been written within the activities of GNCS and GNFM groups of INdAM (Italian National Institute of High Mathematics). G.D. and F.F. thank the Italian Ministry of University and Research (MUR) through the PRIN 2020 project (No. 2020JLWP23) “Integrated Mathematical Approaches to Socio Epidemiological Dynamics”. L.P. acknowledges the support by the Royal Society under the Wolfson Fellowship “Uncertainty quantification, data-driven simulations and learning of multiscale complex systems governed by PDEs” and by MIUR-PRIN 2022 Project (No. 2022KKJP4X), “Advanced numerical methods for time dependent parametric partial differential equations with applications”. The partial support by ICSC – Centro Nazionale di Ricerca in High Performance Computing, Big Data and Quantum Computing, funded by European Union – NextGenerationEU is also acknowledged.

### v2

@@ 5 Conclusion
In this study, we explored the impact of a data augmentation strategy on enhancing the forecasting capabilities of Feed-Forward Neural Networks and Nonlinear Autoregressive Networks in analyzing and predicting the dynamics of the COVID-19 epidemic during the lockdown and post-lockdown phase in Italy and Spain. By integrating real data with synthetic data generated from a generalized SIR model, we demonstrated the effectiveness of neural networks in capturing complex epidemic dynamics. FNNs performed well in interpolating within observed data, while NARs were particularly effective in capturing temporal dependencies, making them suitable for long-term forecasting. The generalized SIR model provided a robust foundation for generating high-resolution synthetic data, enriching the neural network training process. The key outcome of this work is the development of a hybrid methodology that bridges the gap between traditional mathematical modeling and machine learning techniques, such as Physics-Informed Neural Networks, offering a scalable and flexible framework for epidemic modeling and prediction in real-world scenarios.
@@
Acknowledgements
This work has been written within the activities of GNCS and GNFM groups of INdAM (Italian National Institute of High Mathematics). G.D. has been partially funded by the European Union — NextGenerationEU, MUR–PRIN 2022 PNRR Project No. P2022JC95T “Data-driven discovery and control of multi-scale interacting artificial agent systems”. G.D. and F.F. thank the Italian Ministry of University and Research (MUR) through the PRIN 2020 project (No. 2020JLWP23) “Integrated Mathematical Approaches to Socio–Epidemiological Dynamics”. L.P. has been partially funded by the European Union– NextGenerationEU under the program “Future Artificial Intelligence– FAIR” (code PE0000013), MUR PNRR, Project “Advanced MATHematical methods for Artificial Intelligence– MATH4AI”. L.P. acknowledges the support by the Royal Society under the Wolfson Fellowship “Uncertainty quantification, data-driven simulations and learning of multiscale complex systems governed by PDEs” and by MIUR-PRIN 2022 Project (No. 2022KKJP4X), “Advanced numerical methods for time dependent parametric partial differential equations with applications”. The partial support by ICSC – Centro Nazionale di Ricerca in High Performance Computing, Big Data and Quantum Computing, funded by European Union – NextGenerationEU is also acknowledged.